OpenStudy (anonymous):

What is the angular position in radians of the minute hand of a clock at 7:00?

OpenStudy (anonymous):

do you want it with respect to hour hand?????????

OpenStudy (anonymous):

1 hour for 2pi \[\theta = 7 \times 2\pi = 14\pi \] theta = 14pi rad

OpenStudy (anonymous):

@mikyla.simones